### v1

Abstract
The increasing size of the Key-Value (KV) cache during the Large Language Models long-context inference is the main obstacle for its balance between the deployment cost and task accuracy. To reduce the KV cache size in such scenarios, most previous efforts leveraged on the attention weight to evict non-critical cache tokens. But there is a trade-off in those methods, they usually require major modifiation of the inference infrastructure and significant computation overhead. Base on the fact that the Large Lanuage models are autoregresssive models, we propose LagKV, a KV allocation strategy only relying on straight forward comparison among KV themself. It is a totally attention free method which offers easy integration to the main stream inference platform and comparable performance comparing to other complicated KV compression methods. Results on LongBench and PasskeyRetrieval show that, our approach achieves nearly zero loss when the ratio is 2×2$\times$2 × and $\approx 90\%$ of the original model performance for 8×8$\times$8 ×. Especially in the 64-digit passkey retrieval task, our mehod outperforms the attention weight based method $H_{2}O$ over $60\%$ with same compression ratios. Our code is available at https://github.com/AI-Lab-China-Merchants-Bank/LagKV.
1 Introduction
Large Language Models (LLMs) have recently demonstrated remarkable success across diverse text processing tasks, including document retrieval Laban et al. (2023), code generation Gu (2023), and mathematical reasoning (like R1 model DeepSeek-AI et al. (2025)). The Scaling law Kaplan et al. (2020) suggests that larger models generally achieve superior performance. The R1 model further indicates that longer generation sequences with additional ’thinking tokens’ can enhance reasoning capabilities. However, these improvements comes at a significant cost: the growing KV cache size poses a major challenge for efficient LLM inference. Many efforts try to mitigate this challenge.
Most of LLMs are totally relying on Self-Attention mechanism to determine which historical tokens are important in the next token prediction. Therefore, many KV compression approaches are based on it to drop unimportant ones  (Zhang et al., 2024; Xiao et al., 2023; Liu et al., 2024b; Li et al., 2024; Feng et al., 2024). This kind of algorithms keeps a remarkable performance even when the compression ratio is high. However, most of these importance-based token-dropping approaches are incompatible with Flash Attention(FA) because they require computing attention weights to determine token importance. This limitation makes them impractical for deployment, given FA’s critical role in long-context inference.
Another prominent direction in KV cache optimization involves quantization techniques  (Yang et al., 2024; Liu et al., 2024c), which aim to compress the memory footprint of key-value (KV) states by representing them with reduced precision. These methods achieve significant memory savings—often by 4× or more—while preserving model performance through careful error mitigation strategies. Beyond memory efficiency, quantization also reduces the bandwidth overhead of transferring KV cache across devices in distributed inference scenarios, accelerating multi-GPU or memory-bound workloads. However, a critical limitation of pure quantization approaches is that they retain all historical tokens, leaving the computational cost of attention unchanged. For long-context tasks, this means the quadratic complexity of attention persists despite the reduced memory usage.
The simple but with limited performance methods are usually based on the sliding window tokens eviction. Sliding window-based eviction methods—such as those used in Longformer Beltagy et al. (2020), Infinite-LLM Han et al. (2024), and StreamingLLM Xiao et al. (2023)—retain only the initial cache tokens and those within a fixed sliding window, discarding the rest. However, this indiscriminate eviction strategy often leads to a notable degradation in generation quality.
Recent work by  (Liu et al., 2024a; c) addresses the statistical properties of KV states, revealing distinct distribution patterns for keys and values. Their findings suggest that per-channel quantization for keys (which exhibit consistent variance across feature dimensions) and per-token quantization for values (which vary more significantly across sequence positions) yield better fidelity. This observation motivates our key insight: token importance for eviction—traditionally derived from attention weights—can instead be inferred from token- and channel-wise distribution patterns in the KV space. By leveraging these structural properties, we can design a pruning criterion, LagKV, that is both hardware-friendly (compatible with FA) and semantically aware, enabling compute savings alongside memory reduction.
Figure 1: LagKV recursively compression process: partition the KV cache and use the next join chunk as reference to compress the current one. Keep the rest of them as the sliding winow.
2 Methodology
In this section, we formally introduce our KV compression method, LagKV. We begin by looking at the autoregressive process of the LLMs. Inspired by this, we propose a simple yet effective strategy to use the subsequent tokens to compress the previous ones.
2.1 Preliminaries
LLMs’ next token prediction relies on the previous tokens. First, in the prefill stage, the model uses its tokenizer to convert the words to $n$ indices of the embedding metrics $E\in\mathbb{R}^{V\times d}$ of the model and collects the representations to form a input matrix, $X\in\mathbb{R}^{n\times d}$. This matrix is the initial tokens of the first layer of LLM and then each layer will output a same shape matrix as next layer’s input. To depict the operations in each layer, we follow the notation system from  Liu et al. (2023) with $h$ attention heads. For each head $i\in[1,h]$ and head dimension $d_{h}$, we focus on the Query, Key, and Value states, which are converted from tokens by three linear transformation matrices $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}\in\mathbb{R}^{d\times d_{h}}$ sperately:
$Q_{i}=XW_{i}^{Q},K_{i}=XW_{i}^{K},V_{i}=XW_{i}^{V}$ (1)
The output $Y\in\mathbb{R}^{n\times d}$ is computed using the attention weights $A_{i}\in\mathbb{R}^{n\times n}$ and the final output matrix $W^{O}\,\in\mathbb{R}^{d\times d}$:
$Y=Concat_{i\in[1,h]}(A_{i}V_{i})W^{O}\\ \>\text{where}\>A_{i}=\text{softmax}(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{h}}})$ (2)
When the new tokens are generated subsequently in the autoregressive inference, which named as decode stage, the embedding of generated token $x$ is mapped to its respective Query, Key, and Value states for each head, and the previous KV cache is updated accordingly:
$q_{i}=xW_{i}^{Q},k_{i}=xW_{i}^{K},v_{i}=xW_{i}^{V}$ (3)
$K_{i}=Cat[K_{i}:k_{i}],V_{i}=Cat[V_{i}:v_{i}],\\ A_{i}=\text{softmax}(\frac{q_{i}K_{i}^{T}}{\sqrt{d_{h}}})$ (4)
Since $q_{i}\in\mathbb{R}^{1\times d_{h}}$, the computation will be much faster because of the KV cache.
2.2 LagKV
Since the intrinsic property of autoregressive model, the next generated token will not change abruptly from the previous one. As observed in Liu et al. (2024a), the called token-wise locality will show that the tokens in closer proximity have more similar K/V tensor values compared to tokens that are further apart. And also, the StreamingLLM method Xiao et al. (2023) has already proved that the head part and the sliding window of KV cache are very important. Inspired by these, we proposed our LagKV method as:
  • After the prefill is done, start to apply the compression dynamically.
  • Always keep the attention sink with size $S$.
  • Partition the rest KV after the sink part with lag size $L$. If it’s not divisible by $L$, the modulo of it will be added to the sliding window.
  • Recursively compute the KV cache score. Use the next partition as a reference, calculate token-wise max and min from the reference then use max-min to normalize the Key and Value states respectly. After the KV are normalized, calculate the channel-wise standard deviation then softmax. The equations are formally like:
    $min_{i}^{p,K,V}=min_{seq}({K_{i}^{p+1},V_{i}^{p+1}})$ (5)
    $max_{i}^{p,K,V}=max_{seq}({K_{i}^{p+1},V_{i}^{p+1}})$ (6)
    $\bar{K_{i}^{p}},\bar{V_{i}^{p}}=\frac{{K_{i}^{p},V_{i}^{p}}-min_{i}^{p,K,V}}{max_{i}^{p,K,V}-min_{i}^{p,K,V}}$ (7)
    $score(K_{i},V_{i})=Softmax(Std.(\bar{K_{i}},\bar{V_{i}}))$ (8)
    where $p$ denotes the partition index, $i$ represents the head index and $seq$ for the sequence axis. Since the last partition has no reference can be used. It will serve as another part of the sliding window.
  • Sum the scores of Key and Value to get the final score of each token:
    $score_{i}=score(K_{i})+score(V_{i})$ (9)
  • Base on the $score_{i}$, use the Top-k strategy to select tokens in each partition and each head.
The min-max normalization is applied along the sequence dimension, meaning each channel is normalized using statistics from lag-$L$ tokens. Due to token-wise locality, the channel-specific norms of $K_{i}$ and $V_{i}$ are largely eliminated. The resulting normalized representations, $\bar{K_{i}}$ and $\bar{V_{i}}$, retain the original channel-wise variance, allowing the standard deviation to serve as a measure of token importance. The softmax operation then identifies and separates outliers, while the summed scores $score(K_{i})$ and $score(V_{i})$ determine their relative contributions.
As showed in Fig. 1, our method is recursively compressing KV cache in both prefill and decode stages, which is essential for the token-wise locality as mentioned above. It requires relative short distance to keep the similarity among the KV states. Subsequently, another benefit, it also avoids the bias from the long context with length much larger than $L$ and the case when the question is at the end of the prompt Yuan et al. (2024); Tang et al. (2024).
We do not compare the LagKV score to the attention weights here. The attention weights vary on different incoming queries. But our scoring method does not depend on the query states or the tokens after the next joint chunk. It mainly finds the tokens that are not coherent to the next chunk and keep them in the cache. Just like the quantization in KIVI Liu et al. (2024c), we need a rightful mean to find the correct variance and then prune the small ones. However, we use this strategy to evict tokens instead of quantizing them.
To caculate the compresstion ratio, we set the retained token ratio as $r$ in each partition. In the partition chunk, only $rL$ tokens will be kept and others are evicted. Therefore, the compression ratio $C$ for the token sequence length $L_{s}$ not less than $S+2L$ can be expressed as:
$L_{R}=S+rL(Floor(((L_{s}-S)/L)-1)+L+Mod(L_{s}-S,L)$ (10)
$C=1-\frac{L_{R}}{L_{s}}$ (11)
Where $L_{R}$ is the length of the KV cache after compression. For the case $L_{s}\leq S+2L$, the compression ratio is zero.
3 Experiments
3.1 Settings
Base Models. We employ two open-source base models: Llama-3.1-8B-Instruct Grattafiori et al. (2024) and Qwen2.5-7B-Instruct Jiang et al. (2023). These models are main stream LLMs with moderate size and both leverage the GQA Ainslie et al. (2023) technique to reduce the KV cache size.
Datasets. We use the facility in Yuan et al. (2024) to extensive test our method. It mainly contains two benchmarks: LongBench Bai et al. (2023) and Needle-in-a-HaystackTest with Passkey-Retrieval Kamradt (2023); Mohtashami & Jaggi (2023). We only test the 64-digit passkey retrieval task which is much more challenging and use the partial match score in the report.
Parameter Spaces. Across the whole paper, we fix the sink size to $S=16$ and vary the lag size $L$ and ratio $r$. The values of $L$ will be $L=128,512,1024$. The values of $r$ will be 2×2$\times$2 ×, 4×4$\times$4 ×, 6×6$\times$6 × and 8×8$\times$8 × which correspond to $r=0.5,0.25,0.167,0.125$ respectly.
Main results. The main results of this work is Table 1.
Table 1: Performance of LagKV.
Model Method Single. QA Multi. QA Summ. Few-shot Synthetic Code LB Avg. Needle
Llama-3.1-8B-Instruct Baseline 40.71 37.90 28.29 68.49 68.00 58.70 47.44 99.44
L=1024,r=2x 39.42 37.12 27.38 67.71 68.50 58.83 46.74 99.27
L=1024,r=4x 37.06 36.77 26.79 66.96 63.50 58.42 45.54 96.57
L=1024r=,6x 35.74 36.08 26.33 66.33 60.50 57.91 44.65 91.77
L=1024,r=8x 35.49 35.99 25.90 65.21 61.00 57.95 44.31 86.26
L=512,r=2x 39.43 37.45 27.35 67.82 67.50 58.66 46.73 97.02
L=512,r=4x 37.39 36.27 26.19 66.56 62.50 57.86 45.16 85.73
L=512,r=6x 34.95 35.62 25.52 65.87 59.50 58.14 44.11 75.67
L=512,r=8x 34.03 36.12 25.25 64.94 56.00 57.50 43.47 68.25
L=128,r=2x 38.56 36.80 27.20 67.64 68.00 59.27 46.48 92.76
L=128,r=4x 36.66 36.58 25.62 66.78 66.50 57.90 45.28 73.41
L=128,r=6x 34.57 35.41 24.59 63.59 64.00 56.97 43.49 38.48
L=128,r=8x 33.78 34.60 23.91 62.21 61.50 55.68 42.42 25.01
Qwen-2.5-7B-Instruct Baseline 41.62 45.00 26.41 68.91 100.00 63.60 51.53 100.00
L=1024,r=2x 39.80 42.85 26.11 67.66 99.50 63.12 50.33 99.75
L=1024,r=4x 36.92 40.39 24.81 65.91 95.00 61.60 48.15 96.98
L=1024,r=6x 35.77 39.74 24.68 65.28 93.50 61.45 47.52 77.47
L=1024,r=8x 34.60 39.10 24.18 64.74 90.50 61.30 46.73 66.88
L=512,r=2x 38.72 42.79 25.91 67.98 98.50 62.00 49.91 97.07
L=512,r=4x 35.42 39.12 24.49 64.59 94.00 60.16 47.01 75.89
L=512,r=6x 34.00 38.04 23.72 64.31 87.50 58.80 45.69 42.70
L=512,r=8x 32.14 37.83 23.11 63.48 82.50 58.71 44.64 30.00
L=128,r=2x 38.67 42.49 25.69 67.75 99.00 60.64 49.61 65.93
L=128,r=4x 34.47 39.78 24.07 65.13 96.00 58.67 46.91 20.83
L=128,r=6x 32.83 38.15 22.95 62.23 90.50 56.25 44.76 16.18
L=128,r=8x 32.47 37.10 22.20 60.24 88.50 56.10 43.78 15.07
3.2 LongBench
For LongBench dataset, the LagKV method performs very well across different ratios and lag sizes. When $r=2\times$, for different $L$, it shows nearly zero loss compared to the baseline in both models. When $L=1024,r=8\times$, the method still retains approximate $90\%$ of the baseline performance. Since the compression ratio will increase when $L$ decreases, the worse case is $L=128,r=8\times$ for both models but the method maintains at least $85\%$ of the baseline performance.
3.3 Passkey Retrieval
The 64-digit passkey retrieval task is a challenging one for most token eviction strategies. As discussed in Yuan et al. (2024), the most succesful eviction strategy $H_{2}O$ Zhang et al. (2024) performs well in 7-digit task (scoring $100\%$ for all compression ratios) but degrades a lot in the 64-digit one (scoring $35\%$ for 4×4$\times$4 × in Llama-3) since the first token leakage issue. However, our method performs very well when $rL$ is sufficient large enough (scoring $96.57\%$ when $L=1024,r=4\times$). Our recursive compression strategy will not perform well for the setups with small $rL$ due to the fact that digitals usually require more tokens to be represented than the same length words. As shown in Fig. 2, the Qwen model which uses one token for one digit degenerates faster than the Llama model which represents three digits by one token with higher compression ratios. It hints us that we must choose the compression ratio and the lag size carefully in considering the length of the expected content.  A.1 shows all the details of the needle results.
(a) Llama-3.1-8B-Instruct
(b) Qwen-2.5-7B-Instruct
Figure 2: The needle score vs different setups of rL. The horizontal dash-dot line is the baseline for each model. The x-axis is in log scale. We put two vertical lines: $x=64$ (solid blue) and $x=128$ (green dash) for guidelines.
4 Related Works
The $L_{2}$ Norm-Based KV compression  Devoto et al. (2024) is an existing eviction approach that relies solely on KV information to compress the KV cache. This method computes token scores using the negative norm of key states. However, unlike our approach, it does not employ recursive compression during the prefill stage. As discussed in A.2, this method shows limitations in the 64-digit passkey retrieval task when the method is adapted to a recursive framework.
5 Conclusion
In this study, we propose LagKV, an attention-weight-free token eviction method. It achieves comparable performance on long-context tasks while significantly outperforming mainstream eviction strategies in 64-digit passkey retrieval tasks. These results demonstrate that our method maintains robust long-text retrieval capabilities even at high compression ratios.
Unlike existing approaches, LagKV employs a recursive, attention-weight-free strategy in both prefill and decode stages to determine token importance for future processing. It’s indepent from query states and the rest part of the long prompt. Therefore our method offers a novel perspective on LLM mechanisms, shedding light on their inner workings in a fundamentally different way.
Appendix A Appendix
A.1 Detail Rresults of Passkey Retrieval
Here, we present all the Needle-in-a-Haystack results with 64-digit Passkey Retrieval for different setups in Fig.3 and 4.
$L=1024,r=2\times$
$L=1024,r=4\times$
$L=1024,r=6\times$
$L=1024,r=8\times$
$L=512,r=2\times$
$L=512,r=4\times$
$L=512,r=6\times$
$L=512,r=8\times$
$L=128,r=2\times$
$L=128,r=4\times$
$L=128,r=6\times$
$L=128,r=8\times$
Figure 3: The 64-digit Passkey Retrieval of Llama-3.1-8B-Instruct for different setups.
$L=1024,r=2\times$
$L=1024,r=4\times$
$L=1024,r=6\times$
$L=1024,r=8\times$
$L=512,r=2\times$
$L=512,r=4\times$
$L=512,r=6\times$
$L=512,r=8\times$
$L=128,r=2\times$
$L=128,r=4\times$
$L=128,r=6\times$
$L=128,r=8\times$
Figure 4: The 64-digit Passkey Retrieval of Qwen-2.5-7B-Instruct for different setups.
A.2 Variants From LagKV Framework
Here we present two variants from LagKV. Both of them will only change the scoring methods but keep the attention sink and sliding window unchanged. And we only use the 64-digits passkey retrieval task which can easily distinguish eviction strategies as the detector. Among these tests, we keep $S=16,L=1024$ as constant.
The first one is called LocalKV which only skips using the reference from the next joint chunk tokens but replacing the equation Eq. 5 and 6 by the following equations:
$min_{i}^{p,K,V}=min_{seq}({K_{i}^{p},V_{i}^{p}})$ (12)
$max_{i}^{p,K,V}=max_{seq}({K_{i}^{p},V_{i}^{p}})$ (13)
Therefore, the min-max is totally from the local chunk instead of the remote one.
The second one is $L_{2}$ norm from  Devoto et al. (2024). We adapt the low key states norm method into the recursive framework by replacing Eq.9 by:
$score_{i}=-Norm(K_{i})$ (14)
As suggested in their work, we skip the compression of the first two layers in this variant too.
The results of the 64-digits passkey retrieval task is present in Fig. 5. As we can see, the LagKV method is always the best one especially in the high compression ratios. The $L_{2}$ norm method shows very limited performance.
Llama-3.1-8B-Instruct
Qwen-2.5-7B-Instruct
Figure 5: The 64-digit Passkey Retrieval scores of different variants and compression ratios.